At STP, 30cm*3 of H2 gas was burnt in excess oxygen gas. if the final volume of gases in the reaction was 100cm*3., what was the initial volume of the oxygen?


1
Expert's answer
2021-08-23T02:56:37-0400

At STP the volume = 22.4

= 100 × 22.4 = (30 + x) × 22.4

100 = 30+ x

X = 70

= 70cm3